# On the multiplicity of weak solutions for a class of coupled quasilinear elliptic systems

### v3

@@ 1. Introduction
    ($a.1$) $|A_{k}(x,s)|,|D_{s}A_{k}(x,s)|\leq C_{0},$
    ($a.2$) $A_{k}(x,s)\geq\nu>0,$
    for a.e. $x\in\Omega$, for every $s\in\mathbb{R}$,
@@
    ($g.2$) $|s|\geq R\implies 0<\mu G(x,s)\leq\sum_{k=1}^{d}s_{k}G_{s_{k}}(x,s),$
    ($a.3$) $|s|\geq R\implies 0\leq sD_{s}A_{k}(x,s)\leq\gamma A_{k}(x,s)$
    for a.e. $x\in\Omega$.
@@
where $A(x,s)$ is a matrix with $C^{1}$-Carathéodory coefficients.
@@
Several classes of elliptic systems have been investigated in the literature through different analytical approaches. Variational systems—namely those that arise as Euler–Lagrange equations of a suitable functional—can be studied by means of smooth or nonsmooth critical point theory, depending on the regularity of the associated energy functional. Classical examples are given by gradient systems of the form
@@
for which existence and multiplicity of solutions can be obtained through minimization techniques, Mountain Pass arguments, or linking constructions; see, for instance, [15].
@@
More recently, in [8, 9], the authors consider the problem (1) with $d=2$, adopting a different variational framework and assuming a stronger growth condition on the nonlinearities $g_{1}(x,u,v),g_{2}(x,u,v)$.
In the present paper, we assume instead the weaker growth condition ($g.1$), namely the one used in the regular case [3], and we allow $d\geq 2$.
@@ 2.1. Weak slope
We recall some results about the critical point theory of continuous functionals, developed in [12]. In this setting, we consider $X$ a metric space and $f:X\to\mathbb{R}$ a continuous functional.
@@
Let $X$ be a metric space and let $f:X\to\mathds{R}$ be a continuous function. We consider $\sigma\geq 0$ such that there exist $\delta>0$ and a continuous map $\mathscr{H}:B_{\delta}(u)\times[0,\delta]\to X$ such that
(2.1) $d(\mathscr{H}(v,t),v)\leq t,$
@@
Let $X$ be a metric space and let $f:X\to\mathds{R}$ be continuous. We say that $u\in X$ is a (lower) critical point if $|df|(u)=0$. A (lower) critical point is said at the level $c\in\mathbb{R}$ if it is also true that $f(u)=c$.
@@
Furthermore, we will say that $f$ satisfies the $(PS)_{c}$-condition if every $(PS)_{c}$-sequence admits a convergent subsequence in $X$. If the $(PS)_{c}$-condition holds for every $c\in\mathbb{R}$ we will simply write $(PS)$-condition.
@@
  • $\exists\ \rho>0,\alpha>f(0)$ and a subspace $X_{2}\subset X$ of finite codimension such that
    $f\geq\alpha\ on\ \partial B_{\rho}\cap X_{2}$,
  • for every subspace $X_{1}^{(k)}$ of dimension $k$, there exists $R=R^{(k)}>0$ such that $f\leq f(0)$ in $B_{R}^{c}\cap X_{1}^{(k)}$.
If $f$ satisfies the $(PS)_{c}$-condition for every $c\geq\alpha$, then there exists a divergent sequence of critical values, namely there exists a sequence $\{u_{n}\}\subset X$ such that $c_{n}:=f(u_{n})\to+\infty$.
@@ 2.2. Properties of energy functional
where $p_{k}^{\prime}$ is the conjugate exponent of $p_{k}$.
@@
The continuity of the functional follows from the fact that $A_{k}(x,u_{k})$ are continuous in $u_{k}$ and by the Lebesgue’s Theorem. Let, now, $\varphi\in C_{c}^{\infty}(\Omega;\mathbb{R}^{d})$ . Then by Lebesgue’s Theorem, the quantity
@@
$\langle f^{\prime}(u),\varphi\rangle=\limsup_{\begin{subarray}{c}w\to u\\ \ t\to 0^{+}\end{subarray}}\frac{f(w+t\varphi)-f(w)}{t}\geq-|df|(u),\qquad\text{if $\|\varphi\|=1$.}$
@@ 3. $(iii)$
$v_{n,k}=v_{k}\bigg(|v_{k}|^{p_{k}}+\frac{1}{n^{p_{k}}}\bigg)^{-\frac{1}{p_{k}}}\min\bigg\{\bigg(|v_{k}|^{p_{k}}+\frac{1}{n^{p_{k}}}\bigg)^{\frac{1}{p_{k}}}-\frac{1}{n},\bigg(|\tilde{v}_{n,k}|^{p_{k}}+\frac{1}{n^{p_{k}}}\bigg)^{\frac{1}{p_{k}}}-\frac{1}{n}\bigg\}\in W_{0}^{1,p_{k}}\cap L^{\infty}.$
@@
$T_{k}v_{n,k}=\lambda_{n,k}T_{k}v_{k}\geq-\lambda_{n,k}\left[D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}\right]^{-}\geq-\left[D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}v_{k}\right]^{-}.$
@@
$=p_{k}\liminf_{n\to+\infty}\bigg\{\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla v_{n,k}-\langle\omega_{k},v_{n,k}\rangle\bigg\}$
@@ 3. Regularity result
We adapt the procedure in [22], see also [14] and [17]. For every $\gamma,t,L$ such that $L>R+1$, we define
$\tilde{h}_{L,\gamma}(s)=\begin{cases}|s|^{\gamma-1}s,&|s|\leq L,\\ \gamma L^{\gamma-1}s-\operatorname{sign}(s)(1-\gamma)L^{\gamma},&|s|>L,\end{cases}$
and $h_{L,\gamma}(s)$ the function such that
  • $h_{L,\gamma}\equiv 0$ for $|s|\leq R$;
  • $h_{L,\gamma}$ is the line which joins $(R,0)$ with the point $(R+1,(R+1)^{\gamma})$ for $s\in(R,R+1]$ and it is the line which joins (R,0)(-R,0) with $(-R-1,-(R+1)^{\gamma})$ for $s\in[-R-1,-R)$;
  • $h_{L,\gamma}\equiv\tilde{h}_{L,\gamma}$ for $|s|>R+1$.
Since $h_{L,\gamma}$ is differentiable a.e. in $\mathbb{R}$ and its derivative is bounded, we can also define:
$\Phi_{L,t,\gamma}(s)=\int_{0}^{s}|h_{L,\gamma}^{\prime}(r)|^{\frac{t}{\gamma}}\,dr.$
Notice that for every $|s|>L$ we have $h_{L,\gamma}^{\prime}\equiv\gamma L^{\gamma-1}$ and this implies that $\Phi_{L,t,\gamma}(s)=c_{1}(L,\gamma,t)s$. Thus, if $|u_{k}|>L$ we have
$D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}\Phi_{L,t,\gamma}(u_{k})\geq 0$
by ($a.3$) and, in particular, $D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}\Phi_{L,t,\gamma}(u_{k})]^{-}\in L^{1}(\Omega)$. According to Theorem 2.9, $\Phi_{L,t,\gamma}$ is an admissible test function. Now, we start testing (1.1) with $\Phi_{L,\gamma p_{k},\gamma}(u_{k})e_{k}$ where $\{e_{1},\dots,e_{d}\}$ is the canonical basis of $\mathbb{R}^{d}$.
From ($a.3$), we obtain that
$\frac{1}{p_{k}}\int_{\Omega}D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}\Phi_{L,\gamma p_{k},\gamma}(u_{k})\geq 0.$
Indeed, $h_{L,\gamma}^{\prime}(s)$ is a constant for $|s|\leq R+1$ and for $|s|>L$, which implies that $\Phi_{L,\gamma p_{k},\gamma}=\theta s$ for some $\theta\geq 0$ and for every $s\in\mathbb{R}$ such that either $|s|\leq R+1$ or $|s|>L$. Otherwise, $h_{L,\gamma}^{\prime}(s)=\gamma|s|^{\gamma-1}$ and $\Phi_{L,\gamma p_{k},\gamma}(s)=s|s|^{\gamma-1}$. Since
$|s|\geq R\implies sD_{s}A_{k}(x,s)\geq 0,$
we obtain that $D_{s}A_{k}(x,s)\Phi_{L,\gamma p_{k},\gamma}(s)\geq 0$ and the claim follows. Furthermore:
$\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla\Phi_{L,\gamma p_{k},\gamma}(u_{k})=\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}|h_{L,\gamma}^{\prime}(u_{k})|^{p_{k}}.$
Since $h_{L,\gamma}(u_{k})\in W_{0}^{1,p_{k}}(\Omega)$, the Sobolev embedding and (1.1) imply that
$\left(\int_{\Omega}|h_{L,\gamma}(u_{k})|^{q_{k}}\right)^{\frac{p_{k}}{q_{k}}}$ $\leq c\int_{\Omega}|\nabla h_{L,\gamma}(u_{k})|^{p_{k}}=\int_{\Omega}|\nabla u_{k}|^{p_{k}}\cdot|h_{L,\gamma}^{\prime}(u_{k})|^{p_{k}}$
$\leq\frac{1}{\nu}\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}|h_{L,\gamma}^{\prime}(u_{k})|^{p_{k}}$
$\leq C_{1}\int_{\Omega}\left(1+|u_{k}|^{q_{k}-1}+\sum_{j\neq k}|u_{j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}\right)|\Phi_{L,\gamma p_{k},\gamma}(u_{k})|.$
Now, we can follow equations (2.7)-(2.10) in [22].
@@
$\int_{\{|u_{k}|\geq R+1\}}(|u_{k}|^{q_{k}-1}+1)|\Phi_{L,\gamma p_{k},\gamma}(u_{k})|\leq$ $C_{2}\sigma^{q_{k}-1}\left(\int_{\Omega}|h_{L,\gamma}(u_{k})|^{q_{k}}\right)^{\frac{p_{k}}{q_{k}}\frac{\gamma p_{k}+1-p_{k}}{\gamma p_{k}}}$
$+C_{2}\|u_{k}\|_{L^{q_{k}}(\Omega_{\sigma,u_{k}})}^{q_{k}-p_{k}}\left(\int_{\Omega}|h_{L,\gamma}(u_{k})|^{q_{k}}\right)^{\frac{p_{k}}{q_{k}}}.$
We consider the coupling term:
$\int_{\Omega}|u_{j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}|\Phi_{L,\gamma p_{k},\gamma}(u_{k})|$ $\leq\sigma^{q_{j}\frac{q_{k}-1}{q_{k}}}\int_{\Omega}|\Phi_{L,\gamma p_{k},\gamma}(u_{k})|+\int_{\Omega_{\sigma,u_{j}}}|u_{j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}|\Phi_{L,\gamma p_{k},\gamma}(u_{k})|$
$\leq c_{1}(\sigma,R)+\sigma^{q_{j}\frac{q_{k}-1}{q_{k}}}\int_{\{|u_{k}|\geq R+1\}}\Phi_{L,\gamma p_{k},\gamma}(u_{k})|$
$+\int_{\Omega_{\sigma,u_{j}}}|u_{j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}|\Phi_{L,\gamma p_{k},\gamma}(u_{k})|=:c_{1}(\sigma,R)+I_{j,k}.$
@@
$\sum_{j\neq k}I_{j,k}\leq$ $\sum_{j\neq k}\Bigg\{C_{3}\sigma^{q_{j}\frac{q_{k}-1}{q_{k}}}\left(\int_{\Omega}|h_{L,\gamma}(u_{k})|^{q_{k}}\right)^{\frac{p_{k}}{q_{k}}\frac{\gamma p_{k}+1-p_{k}}{\gamma p_{k}}}$
$+C_{3}\|u_{j}\|_{L^{q_{j}}(\Omega_{\sigma,u_{j}})}^{\frac{q_{j}}{q_{k}}(q_{k}-p_{k})}\left[\left(\int_{\Omega}|h_{L,\gamma}(u_{k})|^{q_{k}}\right)^{\frac{p_{k}}{q_{k}}}+\left(\int_{\Omega}\left|h_{L_{1},\gamma}(u_{j})\right|^{q_{j}}\right)^{\frac{p_{k}}{q_{k}}}\right]\Bigg\},$
where $L_{1}:=L^{\frac{q_{k}}{q_{j}}}$. Thus, [22, eq (2.9)] takes the form:
$\int_{\Omega}|h_{L,\gamma}(u_{k})|^{q_{k}}$ $\leq C_{4}+C_{4}\left(\sigma^{q_{k}-1}+\sum_{j\neq k}\sigma^{\frac{q_{j}}{q_{k}}(q_{k}-1)}\right)^{\frac{\gamma q_{k}}{p_{k}-1}}+C_{4}\sum_{j\neq k}\|u_{j}\|_{L^{q_{j}}(\Omega_{\sigma,u_{j}})}^{\frac{q_{j}}{p_{k}}(q_{k}-p_{k})}\int_{\Omega}|h_{L_{1},\gamma}(u_{j})|^{q_{j}}.$
According to [22, Lemma 2.1], we can choose $\sigma\geq\sigma_{1}$ and $C_{5}>0$ independent on $L$ such that:
$\int_{\Omega}|h_{L,\gamma}(u_{j})|^{q_{j}}\leq C_{5}.$
@@
$\xi(s)=\begin{cases}s+R&\text{if $s<-R$}\\ 0&\text{if $|s|\leq R$}\\ s-R&\text{if $s>R$}.\end{cases}$
@@
$\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla\xi(u_{k})=\int_{\{|u_{k}|\geq R\}}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla u_{k},$
$\int_{\Omega}D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}\xi(u_{k})\geq 0.$
@@
$\nu\left(\int_{\Omega}|\xi(u_{k})|^{p_{k}^{*}}\right)^{\frac{p_{k}}{p_{k}^{*}}}$ $\leq\nu C\int_{\{|u_{k}|>R\}}|\nabla u_{k}|^{p_{k}}\leq\nu C\int_{\Omega}|\nabla\xi(u_{k})|^{p_{k}}\leq C\int_{\Omega}g_{k}(x,u)\xi(u_{k})$
@@
$\nu\left(\int_{\Omega}|\xi(u_{k})|^{p_{k}^{*}}\right)^{\frac{p_{k}-1}{p_{k}^{*}}}$ $\leq C\left(\int_{\{|u|>R\}}|g_{k}(x,u)|^{(p_{k}^{*})^{\prime}}\right)^{\frac{1}{(p_{k}^{*})^{\prime}}}$
$\leq C\left(\int_{\{|u|>R\}}|g_{k}(x,u)|^{m}\right)^{\frac{1}{m}}|\{|u_{k}|>R\}|^{\frac{1}{(p_{k}^{*})^{\prime}}-\frac{1}{m}}.$
On the other hand, for every $M>R$ we have that
@@
$|\{|u_{k}|>M\}|\leq\frac{C_{\nu}}{(M-R)^{p_{k}^{*}}}|\{|u_{k}|>R\}|^{\gamma},\quad\gamma=\frac{p_{k}^{*}}{p_{k}-1}\left(\frac{1}{(p_{k}^{*})^{\prime}}-\frac{1}{m}\right)>1.$

### v4

@@ 1. Introduction
    ($a.1$) $|A_{k}(x,s)|,|D_{s}A_{k}(x,s)|\leq C_{0},\qquad$ for a.e. $x\in\Omega$, for every $s\in\mathbb{R}$,
    ($a.2$) $A_{k}(x,s)\geq\nu>0,\qquad$ for a.e. $x\in\Omega$, for every $s\in\mathbb{R}$,
@@
    ($g.2$) $|s|\geq R\implies 0<\mu G(x,s)\leq\sum_{k=1}^{d}s_{k}G_{s_{k}}(x,s),\quad$ for a.e. $x\in\Omega$,
    ($a.3$) $|s_{k}|\geq R\implies 0\leq sD_{s}A_{k}(x,s_{k})\leq\gamma A_{k}(x,s_{k}),\quad$ for a.e. $x\in\Omega$, for every $k=1,\dots,d$.
@@
where $A(x,s)$ is a matrix with $C^{1}$-Carathéodory coefficients, was established in [10, 11, 12].
@@
Several classes of elliptic systems have been investigated in the literature through different analytical approaches. Variational systems—namely those that arise as Euler–Lagrange equations of a suitable functional—can be studied by means of smooth or nonsmooth critical point theory, depending on the regularity of the associated energy functional. Classic examples are given by gradient systems of the form
@@
for which the existence and multiplicity of solutions can be obtained through minimization techniques, Mountain Pass arguments, or linking constructions; see, for instance, [15].
@@
More recently, in [8, 9], the authors consider the problem (1) with $d=2$, adopting a different variational framework and assuming a stronger growth condition on the nonlinearities $g_{1}(x,u),g_{2}(x,u)$.
In the present paper, we assume the weaker growth condition ($g.1$) instead, namely the one used in the regular case [3], and we allow $d\geq 2$.
@@ 2.1. Weak slope
We recall some results about the critical point theory of continuous functionals, developed in [12]. In this setting, we consider $(X,\operatorname{dist})$ a metric space and $f:X\to\mathbb{R}$ a continuous functional.
@@
Let $(X,\operatorname{dist})$ be a metric space and let $f:X\to\mathds{R}$ be a continuous function. We consider $\sigma\geq 0$ such that there exist $\delta>0$ and a continuous map $\mathscr{H}:B_{\delta}(u)\times[0,\delta]\to X$ such that
(2.1) $\operatorname{dist}(\mathscr{H}(v,t),v)\leq t,$
@@
Let $X$ be a metric space and let $f:X\to\mathds{R}$ be continuous. We say that $u\in X$ is a (lower) critical point if $|df|(u)=0$. A (lower) critical point is said to be at the level $c\in\mathbb{R}$ if it is also true that $f(u)=c$.
@@
Furthermore, we will say that $f$ satisfies the $(PS)_{c}$-condition if every $(PS)_{c}$-sequence admits a convergent subsequence in $X$. If the $(PS)_{c}$-condition holds for every $c\in\mathbb{R}$, we will simply write $(PS)$-condition.
@@
  • $\exists\ \rho>0,\alpha>f(0)$ and a subspace $W\subset X$ of finite codimension such that
    $f\geq\alpha\ on\ \partial B_{\rho}\cap W$,
  • for every subspace $V$ of finite dimension , there exists $R=R(V)>0$ such that $f\leq f(0)$ in $B_{R}^{c}\cap V$.
If $f$ satisfies the $(PS)_{c}$-condition for every $c\geq\alpha$, then there exists a divergent sequence of critical values, namely there exists a sequence of critical points $\{u_{n}\}\subset X$ such that $c_{n}:=f(u_{n})\to+\infty$.
@@ 2.2. Properties of energy functional
where $p_{k}^{\prime}$ is the conjugate exponent of $p_{k}$ with $k=1,\dots,d$.
@@
The continuity of the functional follows from the fact that $A_{k}(x,u_{k})$ is continuous in $u_{k}$ and by Lebesgue’s Theorem. Let, now, $\varphi\in C_{c}^{\infty}(\Omega;\mathbb{R}^{d})$ . Then, by Lebesgue’s Theorem, the quantity
@@
$\langle f^{\prime}(u),\varphi\rangle=\limsup_{\begin{subarray}{c}w\to u\\ \ t\to 0^{+}\end{subarray}}\frac{f(w+t\varphi)-f(w)}{t}\geq-\|\varphi\|\cdot|df|(u),\qquad\text{if $\|\varphi\|\leq 1$.}$
@@ 3. $(iii)$
$v_{n,k}=v_{k}\bigg(|v_{k}|^{2}+\frac{1}{n^{2}}\bigg)^{-\frac{1}{2}}\min\bigg\{\bigg(|v_{k}|^{2}+\frac{1}{n^{2}}\bigg)^{\frac{1}{2}}-\frac{1}{n},\bigg(|\tilde{v}_{n,k}|^{2}+\frac{1}{n^{2}}\bigg)^{\frac{1}{2}}-\frac{1}{n}\bigg\}\in W_{0}^{1,p_{k}}\cap L^{\infty}.$
@@
$T_{k}v_{n,k}=\lambda_{n,k}T_{k}v_{k}\geq-\lambda_{n,k}\left[D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}v_{k}\right]^{-}\geq-\left[D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}v_{k}\right]^{-}.$
@@
$=p_{k}\liminf_{n\to+\infty}\bigg\{\langle\omega_{k},v_{n,k}\rangle-\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla v_{n,k}\bigg\}$
@@ 3. Regularity result
We adapt the procedure in [22], see also [14] and [17]. For every ${\tilde{\gamma}}>1$ and $L>1$ such that $L>R+1$, we define
$\tilde{h}_{L,{\tilde{\gamma}}}(s)=\begin{cases}|s|^{{\tilde{\gamma}}-1}s,&|s|\leq L,\\ {\tilde{\gamma}}L^{{\tilde{\gamma}}-1}s+\operatorname{sign}(s)(1-{\tilde{\gamma}})L^{\tilde{\gamma}},&|s|>L,\end{cases}$
and $h_{L,{\tilde{\gamma}}}(s)$ the function such that
  • $h_{L,{\tilde{\gamma}}}\equiv 0$ for $|s|\leq R$;
[…]
  • $h_{L,{\tilde{\gamma}}}\equiv\tilde{h}_{L,{\tilde{\gamma}}}$ for $|s|>R+1$.
Since $h_{L,{\tilde{\gamma}}}$ is differentiable a.e. in $\mathbb{R}$ and its derivative is bounded, for every $t>1$, we can also define:
$\Phi_{L,t,{\tilde{\gamma}}}(s)=\int_{0}^{s}|h_{L,{\tilde{\gamma}}}^{\prime}(r)|^{\frac{t}{{\tilde{\gamma}}}}\,dr.$
Notice that $\frac{\Phi_{L,t,{\tilde{\gamma}}}(s)}{s}\geq 0$ for every $s\neq 0$. Thus,
(3.1) $D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}\Phi_{L,t,{\tilde{\gamma}}}(u_{k})\geq 0$
and, in particular, $D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}\Phi_{L,t,\tilde{\gamma}}(u_{k})]^{-}\in L^{1}(\Omega)$. According to Theorem 2.9, $\Phi_{L,t,\tilde{\gamma}}(u_{k})$ is an admissible test function. Now, we start testing (1.1) with $\Phi_{L,\tilde{\gamma}p_{k},\gamma}(u_{k})e_{k}$ where $\{e_{1},\dots,e_{d}\}$ is the canonical basis of $\mathbb{R}^{d}$.
From (3.1), we obtain that
$\frac{1}{p_{k}}\int_{\Omega}D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}\Phi_{L,{\tilde{\gamma}}p_{k},{\tilde{\gamma}}}(u_{k})\geq 0.$
Furthermore:
(3.2) $\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla\Phi_{L,\tilde{\gamma}p_{k},\tilde{\gamma}}(u_{k})=\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}|h_{L,\tilde{\gamma}}^{\prime}(u_{k})|^{p_{k}}.$
Since $h_{L,\tilde{\gamma}}(u_{k})\in W_{0}^{1,p_{k}}(\Omega)$, the Sobolev embedding, (1.1) and (3.2) imply that
$\left(\int_{\Omega}|h_{L,\tilde{\gamma}}(u_{k})|^{q_{k}}\right)^{\frac{p_{k}}{q_{k}}}$ $\leq c\int_{\Omega}|\nabla h_{L,\tilde{\gamma}}(u_{k})|^{p_{k}}=\int_{\Omega}|\nabla u_{k}|^{p_{k}}\cdot|h_{L,\tilde{\gamma}}^{\prime}(u_{k})|^{p_{k}}$
$\leq\frac{1}{\nu}\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}|h_{L,\tilde{\gamma}}^{\prime}(u_{k})|^{p_{k}}$
$\leq C_{1}\int_{\Omega}\left(1+|u_{k}|^{q_{k}-1}+\sum_{j\neq k}|u_{j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}\right)|\Phi_{L,\tilde{\gamma}p_{k},\tilde{\gamma}}(u_{k})|,$
for some constants $c,C_{1}>0$ independent of $L$. Now, we can follow equations (2.7)-(2.10) in [22].
@@
$\int_{\{|u_{k}|\geq R+1\}}(|u_{k}|^{q_{k}-1}+1)|\Phi_{L,\tilde{\gamma}p_{k},\tilde{\gamma}}(u_{k})|\leq$ $C_{2}\sigma^{q_{k}-1}\left(\int_{\Omega}|h_{L,\tilde{\gamma}}(u_{k})|^{q_{k}}\right)^{\frac{p_{k}}{q_{k}}\frac{\tilde{\gamma}p_{k}+1-p_{k}}{\tilde{\gamma}p_{k}}}$
$+C_{2}\|u_{k}\|_{L^{q_{k}}(\Omega_{\sigma,u_{k}})}^{q_{k}-p_{k}}\left(\int_{\Omega}|h_{L,\tilde{\gamma}}(u_{k})|^{q_{k}}\right)^{\frac{p_{k}}{q_{k}}},$
for some $C_{2}>0$ independent of $L$. We consider the coupling term:
$\int_{\Omega}|u_{j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}|\Phi_{L,\tilde{\gamma}p_{k},\tilde{\gamma}}(u_{k})|$ $\leq\sigma^{q_{j}\frac{q_{k}-1}{q_{k}}}\int_{\Omega}|\Phi_{L,\tilde{\gamma}p_{k},\tilde{\gamma}}(u_{k})|+\int_{\Omega_{\sigma,u_{j}}}|u_{j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}|\Phi_{L,\tilde{\gamma}p_{k},\tilde{\gamma}}(u_{k})|$
$\leq c_{1}(\sigma,R)+\sigma^{q_{j}\frac{q_{k}-1}{q_{k}}}\int_{\{|u_{k}|\geq R+1\}}|\Phi_{L,\tilde{\gamma}p_{k},\tilde{\gamma}}(u_{k})|$
$+\int_{\Omega_{\sigma,u_{j}}}|u_{j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}|\Phi_{L,\tilde{\gamma}p_{k},\tilde{\gamma}}(u_{k})|=:c_{1}(\sigma,R)+I_{j,k},$
where $c_{1}(\sigma,R)>0$ is independent of $L$.
@@
$\sum_{j\neq k}I_{j,k}\leq$ $\sum_{j\neq k}\Bigg\{C_{3}\sigma^{q_{j}\frac{q_{k}-1}{q_{k}}}\left(\int_{\Omega}|h_{L,\tilde{\gamma}}(u_{k})|^{q_{k}}\right)^{\frac{p_{k}}{q_{k}}\frac{\tilde{\gamma}p_{k}+1-p_{k}}{\tilde{\gamma}p_{k}}}$
$+C_{3}\|u_{j}\|_{L^{q_{j}}(\Omega_{\sigma,u_{j}})}^{\frac{q_{j}}{q_{k}}(q_{k}-p_{k})}\left[\left(\int_{\Omega}|h_{L,\tilde{\gamma}}(u_{k})|^{q_{k}}\right)^{\frac{p_{k}}{q_{k}}}+\left(\int_{\Omega}\left|h_{L_{1},\tilde{\gamma}}(u_{j})\right|^{q_{j}}\right)^{\frac{p_{k}}{q_{k}}}\right]\Bigg\},$
where $L_{1}:=L^{\frac{q_{k}}{q_{j}}}$ and $C_{3}>0$ is independent of $L$. Thus, [22, eq (2.9)] takes the form:
$\int_{\Omega}|h_{L,\tilde{\gamma}}(u_{k})|^{q_{k}}$ $\leq C_{4}+C_{4}\left(\sigma^{q_{k}-1}+\sum_{j\neq k}\sigma^{\frac{q_{j}}{q_{k}}(q_{k}-1)}\right)^{\frac{\tilde{\gamma}q_{k}}{p_{k}-1}}+C_{4}\sum_{j\neq k}\|u_{j}\|_{L^{q_{j}}(\Omega_{\sigma,u_{j}})}^{\frac{q_{j}}{p_{k}}(q_{k}-p_{k})}\int_{\Omega}|h_{L_{1},\tilde{\gamma}}(u_{j})|^{q_{j}},$
with $C_{4}>0$ independent of $L$. According to [22, Lemma 2.1], we have
$\|u_{j}\|_{L^{q_{j}}(\Omega_{\sigma,u_{j}})}^{\frac{q_{j}}{p_{k}}(q_{k}-p_{k})}\to 0\qquad\text{as $\sigma\to+\infty$},\qquad\text{ for every $j=1,\dots,d$},$
and we can choose $\sigma\geq\sigma_{1}$ and $C_{5}>0$ independent on $L$ such that:
$\int_{\Omega}|h_{L,\tilde{\gamma}}(u_{j})|^{q_{j}}\leq\sum_{k=1}^{d}\int_{\Omega}|h_{L,\tilde{\gamma}}(u_{k})|^{q_{k}}\leq C_{5},\qquad\text{for every $j=1,\dots,d$}.$
@@
$\xi(s)=\begin{cases}0&\text{if $|s|\leq R$}\\ \left(1-\frac{R}{|s|}\right)s&\text{if $|s|>R$}.\end{cases}$
@@
$\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla\xi(u_{k})=\int_{\{|u_{k}|\geq R\}}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla u_{k}.$
According to ($a.3$), we have
$\int_{\Omega}D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}\xi(u_{k})=\int_{\{|u_{k}|\geq R\}}\left(1-\frac{R}{|u_{k}|}\right)D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}u_{k}\geq 0.$
@@
$\nu\left(\int_{\Omega}|\xi(u_{k})|^{p_{k}^{*}}\right)^{\frac{p_{k}}{p_{k}^{*}}}$ $\leq\nu C\int_{\{|u_{k}|>R\}}|\nabla u_{k}|^{p_{k}}\leq C\int_{\Omega}A_{k}(x,u_{k})|\nabla\xi(u_{k})|^{p_{k}}\leq C\int_{\Omega}g_{k}(x,u)\xi(u_{k})$
@@
$\nu\left(\int_{\Omega}|\xi(u_{k})|^{p_{k}^{*}}\right)^{\frac{p_{k}-1}{p_{k}^{*}}}$ $\leq C\left(\int_{\{|u_{k}|>R\}}|g_{k}(x,u)|^{(p_{k}^{*})^{\prime}}\right)^{\frac{1}{(p_{k}^{*})^{\prime}}}$
$\leq C\left(\int_{\{|u_{k}|>R\}}|g_{k}(x,u)|^{m}\right)^{\frac{1}{m}}|\{|u_{k}|>R\}|^{\frac{1}{(p_{k}^{*})^{\prime}}-\frac{1}{m}}.$
On the other hand, for every $M>R$, we have that
@@
$|\{|u_{k}|>M\}|\leq\frac{C_{\nu}}{(M-R)^{p_{k}^{*}}}|\{|u_{k}|>R\}|^{\theta},\quad\theta=\frac{p_{k}^{*}}{p_{k}-1}\left(\frac{1}{(p_{k}^{*})^{\prime}}-\frac{1}{m}\right)>1.$